OpenStudy (anonymous):

Help I will give medal Find an equation for the line with y-intercept 3 that is perpendicular to the line 3y = 2x-4. A. 2y = 6-3x B. 2y = 3x + 6 C. 3y = 9-2x D. 3y = 2x + 9

OpenStudy (anonymous):

2y=6-3x.

OpenStudy (anonymous):

itas a long explanation... if you want i can explain it...

OpenStudy (anonymous):

How did you get it

OpenStudy (anonymous):

Remember if a line has the y-intercept 3, then it goes through the point (0,3)

OpenStudy (anonymous):

two perpendicular lines are negative reciprocals of each other. So if you're given the slope of 2/3 you can find the perpendicular slope by this formula: m (underexponent) P=-1/m

OpenStudy (anonymous):

So plug in the given slope to find the perpendicular slope one over 2/3 (-1/1)*(3/2) then multiply comin out with the anwser of (-3/2) now we know the slope of the unknown line is -3/2 (its the negative reciprocal of 2/3 from the line

OpenStudy (anonymous):

y=(2/3)*x-4 this is what you would plug in y-y(underexponent) 1=m(x-x (underecponent)1 X underexponent 1 Yunderexponent 1

OpenStudy (anonymous):

Then Plug in the following.... m =-3/2 X (underexponent) 1 =0 Y(unerexponent)1= 3 this is what its supposed to look like y-3=(-3/2) *( X-0 )

OpenStudy (anonymous):

multily the fractions and then you come out with the improper fractions and solve

OpenStudy (anonymous):

coming out with 2y = 6-3x i hope this helped.. its eaiser if i can help one on one you know. may i please habe a metal now? :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem thanks for the metal

OpenStudy (anonymous):

Your welcome